# The metric theory of small gaps for a sequence of real numbers

### v1

@@ Abstract.
Let $(a_{n})_{n\geq 1}$ be a sequence of distinct positive integers. The metric theory of minimal gaps for the sequence $\{\alpha a_{n}\text{ mod }1,1\leq n\leq N\}$ as $N\to\infty$ was initiated by Rudnick [Rudnick], who established that the minimal gap admits an asymptotic upper bound expressible in terms of the additive energy of $\{a_{1},\ldots,a_{N}\}$ for almost every $\alpha$. Later, Aistleitner, El-Baz, and Munsch [ABM] demonstrated that the metric theory of minimal gaps for such sequences is governed not by the additive energy, but by the cardinality of the difference set of $\{a_{1},\ldots,a_{N}\}$. They established a sharp convergence test for the typical asymptotic order of the minimal gap and proved general upper and lower bounds that are readily applicable. A key element of their proof relies on the resolution of the Duffin–Schaeffer conjecture by Koukoulopoulos and Maynard. In this article, we generalise several results from [ABM] on integer sequences to the case of real sequences. While an upper bound for $\delta_{\min}^{\alpha}(N)$ remains elusive, we obtain one for its floored counterpart $\lfloor\delta^{\alpha}_{\min}\rfloor(N)$ (both defined in Section 1) for real sequences $(a_{n})_{n\geq 1}$ of distinct numbers. Our theorems recover [ABM, Theorems 1–3], as well as the result from [ABM, Section 4.3]. Furthermore, we establish lower bounds for the minimal gaps of well-spaced sequences and, more generally, of a broader family that contains them.
@@ 1. Introduction
The theory of uniform distribution modulo one is a fundamental area at the intersection of number theory, ergodic theory, and metric Diophantine approximation. A central problem within this field is to understand the local statistics of sequences $(\alpha a_{n})_{n\geq 1}$, where $\alpha$ is a real parameter and $(a_{n})_{n\geq 1}$ is a given sequence of distinct real numbers. A key statistic capturing the finest scale behaviour is the minimal gap, defined for $N$ terms as
@@
where $\|x\|=\displaystyle\min_{k\in\mathbb{Z}}|x-k|$. The behaviour of $\delta_{\min}^{\alpha}(N)$ as $N\to\infty$ for generic $\alpha$ (that is, for almost all $\alpha$ with respect to the Lebesgue measure) reveals deep arithmetic properties of the sequence $(a_{n})_{n\geq 1}$. As direct computation of the minimal gap $\delta_{\min}^{\alpha}(N)$ for fixed $\alpha$ is typically difficult, we analyse its metric properties, focusing on the almost-sure behavior and distribution.
The expected behaviour of the minimal gap serves as a powerful tool for distinguishing random from deterministic sequences. The minimal gap for a random sequence is always at most $1/N$, as $1/N$ is the average spacing. For a sequence of $N$ independent, uniformly distributed random points on the unit interval, the minimal gap is almost surely of size about $1/N^{2}$, a hallmark of Poissonian statistics where points exhibit no correlations. This random benchmark provides a crucial baseline: a deterministic sequence is considered to exhibit Poissonian behaviour if its minimal gap also decays as $1/N^{2}$ as in the random case.
In contrast to the random case, more structured sequences, such as $(\sqrt{n})_{n\notin\square}$, is more rigidly spaced, resulting in a larger minimal gap that decays more slowly (see [Regavim]), thereby signifying a clear departure from random behaviour. It is known that the minimal gap in this case is of order $2/N^{3/2}$, which is much larger than the random scaling of $1/N^{2}$. Since polynomial sequences $(n^{d})_{n\geq 1}$ for $d\geq 2$ and lacunary sequences have nearly minimal additive energy (close to $N^{2}$), the minimal gap $\delta_{\min}^{\alpha}(N)$ is of order $1/N^{2}$ up to sub-power factors. This follows from the following results of Rudnick [Rudnick, Theorems 1–2].
Let $(a_{n})_{n\geq 1}$ be a sequence of distinct integers and its additive energy is nearly minimal, that is, $\ll N^{2+o(1)}$. Then, for almost all $\alpha$, any $\eta>0$ , and for all sufficiently large $N$, we have
@@
The landmark work of Aistleitner, El-Baz, and Munsch [ABM] established that the metric theory of minimal gaps for integer sequences is not controlled by the additive energy, but rather by the cardinality of the difference set $\{a_{i}-a_{j}:1\leq i\neq j\leq N\}$, providing a comprehensive framework that connects the almost-sure minimal gap of $(\alpha a_{n})_{n\geq 1}$ directly to its difference set structure. This article significantly extends their line of research to sequences $(a_{n})_{n\geq 1}$ of arbitrary distinct positive real numbers. We introduce and study several natural variants of the minimal gap $\delta_{\min}^{\alpha}(N)$, namely, $\lfloor\delta_{\min}^{\alpha}\rfloor(N)$, $\widetilde{\delta}_{\min}^{\alpha}(N)$, and $\widehat{\delta}_{\min}^{\alpha}(N)$ (defined in Sections 1 and 3). A primary aim is to understand the relationships between these quantities and to determine their typical asymptotic order for almost every $\alpha$.
Our main results are multifaceted. First, we establish general lower bound and weak upper bound for $\delta_{\min}^{\alpha}(N)$ under the mild assumption that the sequence $(a_{n})_{n\geq 1}$ has a positive uniform lower bound on its gaps, $\inf_{m\neq n}|a_{m}-a_{n}|\geq c>0$. These bounds are expressed in terms of $D_{N}$, defined as the number of positive distinct differences among $\{a_{1},\ldots,a_{N}\}$, and feature the precise logarithmic factors endemic to such metric problems.
@@
The investigation of sequences modulo one sits at the rich confluence of number theory, ergodic theory, and probability. A sequence $(a_{n})_{n\geq 1}$ in $[0,1)$ is considered uniformly distributed if, in the limit, the proportion of its terms falling into any subinterval is equal to the interval’s length. This “global” equidistribution, a concept firmly established since the pioneering work of Weyl [Weyl1916], is elegantly quantified by discrepancy theory, which measures the maximum deviation from this ideal distribution [MichaelTichy, kuipers1974uniform].
However, global uniformity often masks a more complex microscopic reality. Two sequences can be perfectly uniformly distributed yet exhibit radically different local behaviours. The true arithmetic character of a sequence is frequently revealed not by its global distribution, but by its local statistics—the fine-scale patterns such as the distribution of gaps between consecutive points and the correlation between the spacings of multiple points. The study of these statistics, including pair correlation and nearest-neighbor gap distributions, has proven to be a remarkably sensitive probe. It forges deep and often unexpected connections to diverse fields such as random matrix theory, quantum chaos, and metric Diophantine approximation, providing a unifying language for describing randomness and structure in deterministic systems.
While the local statistics of integer sequences have been extensively studied (see [ABM2021, Aistleitner-Larcher-Lewko, Bloom-Chow, Rudnick, Rudnick-Sarnak]), the natural and vast generalisation to sequences of arbitrary real numbers has remained largely unexplored. This gap is significant, as many sequences of central importance in analysis and mathematical physics—such as lacunary sequences, values of smooth functions, and geometric progressions with non-integer ratios—reside firmly within this real-valued realm. A comprehensive theory for real sequences is therefore not merely a technical extension but a necessary step for a complete understanding.
This article is motivated by the goal of establishing such a theory for the minimal gap statistic. We seek to generalise the seminal findings of Aistleitner, El-Baz, and Munsch [ABM] for integer sequences, which tied the minimal gap to the cardinality of the difference set, into the domain of real sequences. By introducing and analysing new, natural variants of the minimal gap tailored for real numbers, we aim to uncover the fundamental principles that govern their fine-scale distribution. Our work provides a unified framework that not only recovers classical results but also opens the door to analysing a much broader and richer class of sequences, thereby deepening our understanding of the interplay between arithmetic structure and random-like behavior.
@@ 1.1. Organisation of the article
Section 3 is dedicated to the study of minimal gap statistics for sequences of distinct positive real numbers. We begin by defining four key metrics: $\delta_{\min}^{\alpha}(N)$, $\widetilde{\delta}_{\min}^{\alpha}(N)$, $\widehat{\delta}_{\min}^{\alpha}(N)$, and $\lfloor\delta_{\min}^{\alpha}\rfloor(N)$. We then establish the fundamental chain of inequalities $\delta_{\min}^{\alpha}(N)\leq\widetilde{\delta}_{\min}^{\alpha}(N)\leq\widehat{\delta}_{\min}^{\alpha}(N)$. The core of this section is the proof of Theorem 1, which provides nearly optimal lower and upper bounds for $\delta_{\min}^{\alpha}(N)$ for sequences with a uniform positive separation $\inf_{m\neq n}|a_{m}-a_{n}|\geq c>0$. The proof employs a careful application of the Borel–Cantelli lemma. We also survey relevant prior results, particularly those of Aistleitner, El-Baz, and Munsch for integer sequences (see [ABM]).
@@
In Section 5, we present a complete zero-one law characterisation for $\lfloor\delta_{\min}^{\alpha}\rfloor(N)\leq\eta(N)$ to occur infinitely often in Theorem 13. The proofs of Lemma 50, Theorem 9, and Theorem 13 build upon corresponding arguments in [ABM] for integer sequences, incorporating key modifications developed in previous lemmas to adapt to our setting. We provide detailed proofs for completeness
@@ 2. Main Results
Before stating the main results, we collect the key notations and definitions used throughout the article.
@@ 2.1. Notation and definitions
  • $\|x\|$: The distance from $x$ to the nearest integer.
@@
  • $\log_{k}x$: The $k$-th iterated logarithm, defined recursively by $\log_{1}x=\log x$ and $\log_{k}x=\log(\log_{k-1}x)$ for $k\geq 2$. Thus $1/\log n$, $\log_{2}n$, $\log_{3}n$, and so on are positive and well-defined for all integers $n\geq 1$.
  • $f(x)\ll g(x)$: $|f(x)|\leq C|g(x)|$ for some constant $C>0$ and all $x$ in a specified domain. Equivalently, $f(x)=O(g(x))$. We write $g(x)\gg f(x)$ to mean $f(x)\ll g(x)$.
@@
  • $\mathcal{N}(k)$: The first index $M$ at which the positive integer $k$ appears in the floored positive difference set, defined by
@@
  • Almost all $\alpha$: A property that holds for all $\alpha$ in a subset of $[0,1]$ with Lebesgue measure 1.
@@ 2.2. General lower and upper bounds
Our first main result provides nearly optimal lower and upper bounds for the minimal gap $\delta_{\min}^{\alpha}(N)$ for sequences with a positive uniform separation. The bounds are given in terms of the quantity $D_{N}$, the number of positive distinct differences among the first $N$ terms.
@@
In particular, the condition $\inf_{m\neq n}|a_{m}-a_{n}|\geq 1$ is automatically satisfied for any sequence $(a_{n})_{n\geq 1}$ of distinct positive integers. Therefore, we recover the second part of [ABM, Theorem 1].
@@ 2.5. A zero-one law for the minimal gap
Let $\left(a_{n}\right)_{n\geq 1}$ be a sequence of positive real numbers such that $\inf_{1\leq m\neq n\leq N}|a_{m}-a_{n}|\geq 1$. Let $(\eta(N))_{N\geq 1}$ be a sequence of non-negative reals. Let $\mathcal{A}$ denote the set of those $\alpha\in[0,1]$ for which $\lfloor\delta_{\min}^{\alpha}\rfloor(N)\leq\eta(N)$ holds for infinitely many $N$. Then, we have $\lambda(\mathcal{A})=0$ or $\lambda(\mathcal{A})=1$, according to whether the series
@@
Theorem 13 provides a sharp criterion reducing the problem of determining the almost sure behavior of the minimal gap of $(\alpha a_{n})_{n\geq 1}$ to the arithmetic structure of $\lfloor(A_{N}-A_{N})^{+}\rfloor$, showing this floored difference set alone determines the typical asymptotic order, at least in principle.
@@ 2.6. Applications and examples
  • Lacunary sequences: Sequences satisfying $\frac{a_{n+1}}{a_{n}}\geq D>1$, such as $a_{n}=e^{n}$ or $\pi^{n}$.
@@
Now taking minimum over all $m^{\prime}\neq n^{\prime}\leq N$, we obtain $\widetilde{\delta}_{\min}^{\alpha}(N)\leq\widehat{\delta}_{\min}^{\alpha}(N)$.
@@
Now taking minimum over all $m^{\prime}\neq n^{\prime}\leq N$, we obtain $\delta_{\min}^{\alpha}(N)\leq\widetilde{\delta}_{\min}^{\alpha}(N)$. Therefore, for all $N\in\mathbb{N}$, we obtain

### v2

@@ Abstract.
Let $(a_{n})_{n\geq 1}$ be a sequence of distinct positive integers. The metric theory of minimal gaps for the sequence $\{\alpha a_{n}\text{ mod }1,1\leq n\leq N\}$ as $N\to\infty$ was initiated by Rudnick [Rudnick], who established that the minimal gap admits an asymptotic upper bound expressible in terms of the additive energy of $\{a_{1},\ldots,a_{N}\}$ for almost every $\alpha$. Later, Aistleitner, El-Baz, and Munsch [ABM] demonstrated that the metric theory of minimal gaps for such sequences is not governed by the additive energy, but by the cardinality of the difference set of $\{a_{1},\ldots,a_{N}\}$. They established a sharp convergence test for the typical asymptotic order of the minimal gap and proved general upper and lower bounds that are readily applicable. A key element of their proof relies on the resolution of the Duffin–Schaeffer conjecture by Koukoulopoulos and Maynard. In this article, we generalise several results from [ABM] on integer sequences to the case of real sequences. Although an upper bound for $\delta_{\min}^{\alpha}(N)$ remains elusive, we obtain one for its floored counterpart $\lfloor\delta^{\alpha}_{\min}\rfloor(N)$ (both defined in Section 1) for real sequences $(a_{n})_{n\geq 1}$ of distinct numbers. Our theorems recover [ABM, Theorems 1–3], as well as the result from [ABM, Section 4.3]. Furthermore, we establish lower bounds for the minimal gaps of well-spaced sequences and, more generally, of a broader family that contains them.
@@ 1. Introduction
The theory of uniform distribution modulo one is a fundamental area at the intersection of number theory, ergodic theory, and metric Diophantine approximation. A central problem within this field is to understand the local statistics of the sequences $(\alpha a_{n})_{n\geq 1}$, where $\alpha$ is a real parameter and $(a_{n})_{n\geq 1}$ is a given sequence of distinct real numbers. A key statistic capturing the finest scale behaviour is the minimal gap, defined for $N$ terms as
@@
where $\|x\|=\displaystyle\min_{k\in\mathbb{Z}}|x-k|$. The behaviour of $\delta_{\min}^{\alpha}(N)$ as $N\to\infty$ for generic $\alpha$ (that is, for almost all $\alpha$ with respect to the Lebesgue measure) reveals the deep arithmetic properties of the sequence $(a_{n})_{n\geq 1}$. As direct computation of the minimal gap $\delta_{\min}^{\alpha}(N)$ for fixed $\alpha$ is typically difficult, we analyse its metric properties, focusing on the almost-sure behaviour and distribution.
The expected behaviour of the minimal gap serves as a powerful tool for distinguishing random from deterministic sequences. The minimal gap for a random sequence is always at most $1/N$, as $1/N$ is the average spacing. For a sequence of $N$ independent and uniformly distributed random points in the unit interval, the minimal gap is almost surely of size about $1/N^{2}$, a hallmark of Poissonian statistics where points exhibit no correlations. This random benchmark provides a crucial baseline: a deterministic sequence is considered to exhibit Poissonian behaviour if its minimal gap also decays as $1/N^{2}$ as in the random case.
In contrast to the random case, more structured sequences, such as $(\sqrt{n})_{n\notin\square}$, are more rigidly spaced, resulting in a larger minimal gap that decays more slowly (see [Regavim]), thus signifying a clear departure from random behaviour. The minimal gap in this case is known to be of the order $2/N^{3/2}$, which is much larger than the random scaling of $1/N^{2}$. Since polynomial sequences $(n^{d})_{n\geq 1}$ for $d\geq 2$ and lacunary sequences have nearly minimal additive energy (close to $N^{2}$), the minimal gap $\delta_{\min}^{\alpha}(N)$ is of order $1/N^{2}$ up to sub-power factors. This follows from the following results of Rudnick [Rudnick, Theorems 1–2].
Let $(a_{n})_{n\geq 1}$ be a sequence of distinct integers, and let its additive energy be nearly minimal, that is, $\ll N^{2+o(1)}$. Then, for almost all $\alpha$, any $\eta>0$ , and for all sufficiently large $N$, we have
@@
The landmark work of Aistleitner, El-Baz, and Munsch [ABM] established that the metric theory of minimal gaps for integer sequences is not controlled by the additive energy, but rather by the cardinality of the difference set $\{a_{i}-a_{j}:1\leq i\neq j\leq N\}$, providing a comprehensive framework that connects the almost-sure minimal gap of $(\alpha a_{n})_{n\geq 1}$ directly to its difference set structure. This article significantly extends their line of research to sequences $(a_{n})_{n\geq 1}$ of arbitrary distinct positive real numbers. We introduce and study several natural variants of the minimal gap $\delta_{\min}^{\alpha}(N)$, namely $\lfloor\delta_{\min}^{\alpha}\rfloor(N)$, $\widetilde{\delta}_{\min}^{\alpha}(N)$, and $\widehat{\delta}_{\min}^{\alpha}(N)$ (defined in Sections 1 and 3). A primary aim is to understand the relationships between these quantities and to determine their typical asymptotic order for almost all $\alpha$.
Our main results are multifaceted. First, we establish a general lower bound and a weak upper bound for $\delta_{\min}^{\alpha}(N)$ under the mild assumption that the sequence $(a_{n})_{n\geq 1}$ has a uniform positive lower bound on its gaps, $\inf_{m\neq n}|a_{m}-a_{n}|\geq c>0$. These bounds are expressed in terms of $D_{N}$, defined as the number of distinct positive differences between $\{a_{1},\ldots,a_{N}\}$, and feature precise logarithmic factors endemic to such metric problems.
@@
The investigation of sequences modulo one sits at the rich confluence of number theory, ergodic theory, and probability. A sequence $(a_{n})_{n\geq 1}$ in $[0,1)$ is considered uniformly distributed if, in the limit, the proportion of its terms that fall into any subinterval is equal to the length of the interval. This “global” equidistribution, a concept firmly established since the pioneering work of Weyl [Weyl1916], is elegantly quantified by discrepancy theory, which measures the maximum deviation from this ideal distribution [MichaelTichy, kuipers1974uniform].
However, global uniformity often masks a more complex microscopic reality. Two sequences can be perfectly uniformly distributed, yet exhibit radically different local behaviours. The true arithmetic character of a sequence is frequently revealed not by its global distribution but by its local statistics, the fine-scale patterns such as the distribution of gaps between consecutive points and the correlation between the spacings of multiple points. The study of these statistics, including pair correlation and nearest-neighbour gap distributions, has proven to be a remarkably sensitive probe. It forges deep and often unexpected connections to diverse fields such as random matrix theory, quantum chaos, and the metric Diophantine approximation, providing a unifying language for describing randomness and structure in deterministic systems.
Although local statistics of integer sequences have been extensively studied (see [ABM2021, Aistleitner-Larcher-Lewko, Bloom-Chow, Rudnick, Rudnick-Sarnak]), with important work on the distribution of spacings and pair correlation for sequences like $(n^{2}\alpha)_{n\geq 1}$ [RudnickSarnakAlexandru2001] and other lacunary sequences [RudnickAlexandru1999, RudnickAlexandru2002], the natural and vast generalisation to sequences of arbitrary real numbers has remained largely unexplored. This gap is significant, as many sequences of central importance in analysis and mathematical physics—such as lacunary sequences, values of smooth functions, and geometric progressions with non-integer ratios—reside firmly within this real-valued realm. A comprehensive theory for real sequences is, therefore, not merely a technical extension but a necessary step for a complete understanding.
This article is motivated by the goal of establishing such a theory for the minimal gap statistic. We seek to generalise the seminal findings of Aistleitner, El-Baz, and Munsch [ABM] for integer sequences, which tied the minimal gap to the cardinality of the difference set, into the domain of real sequences. By introducing and analysing new, natural variants of the minimal gap tailored for real numbers, we aim to uncover the fundamental principles that govern their fine-scale distribution. Our work provides a unified framework that not only recovers classical results but also opens the door to analysing a much broader and richer class of sequences, thereby deepening our understanding of the interplay between arithmetic structure and random-like behaviour.
@@ 1.1. Organisation of the article
Section 3 is dedicated to the study of minimal gap statistics for sequences of distinct positive real numbers. We begin by defining four key metrics: $\delta_{\min}^{\alpha}(N)$, $\widetilde{\delta}_{\min}^{\alpha}(N)$, $\widehat{\delta}_{\min}^{\alpha}(N)$, and $\lfloor\delta_{\min}^{\alpha}\rfloor(N)$. We then establish the fundamental chain of inequalities $\delta_{\min}^{\alpha}(N)\leq\widetilde{\delta}_{\min}^{\alpha}(N)\leq\widehat{\delta}_{\min}^{\alpha}(N)$. The core of this section is the proof of Theorem 1, which provides nearly optimal lower and upper bounds for $\delta_{\min}^{\alpha}(N)$ for sequences with uniform positive separation $\inf_{m\neq n}|a_{m}-a_{n}|\geq c>0$. The proof employs a careful application of the Borel–Cantelli lemma. We also survey relevant prior results, particularly those of Aistleitner, El-Baz, and Munsch for integer sequences (see [ABM]).
@@
In Section 5, we present a complete zero-one law characterisation for $\lfloor\delta_{\min}^{\alpha}\rfloor(N)\leq\eta(N)$ to occur infinitely often in Theorem 13. The proofs of Lemma 49, Theorem 9, and Theorem 13 build upon the corresponding arguments in [ABM] for integer sequences, incorporating key modifications developed in previous lemmas to adapt to our setting. We provide detailed proofs for completeness
@@ 2. Main Results
Before stating the main results, we collect the key notation and definitions used throughout the article.
@@ 2.1. Notation and definitions
  • $\|x\|$: The distance from $x$ to the nearest integer. For $x\in\mathbb{R}$, $\|x\|=\min(\{x\},1-\{x\})$.
@@
  • $\log_{k}x$: The $k$-th iterated logarithm, defined recursively by $\log_{1}x=\log x$ and $\log_{k}x=\log(\log_{k-1}x)$ for $k\geq 2$. Thus, $1/\log n$, $\log_{2}n$, $\log_{3}n$, etc. are positive and well-defined for all integers $n\geq 1$.
  • $f(x)\ll g(x)$: $|f(x)|\leq C|g(x)|$ for some constant $C>0$ and for all $x$ in a specified domain. Equivalently, $f(x)=O(g(x))$. We write $g(x)\gg f(x)$ to mean $f(x)\ll g(x)$.
@@
  • $\mathcal{N}(k)$: The first index $M$ at which the positive integer $k$ appears in the set of floored positive differences, defined by
@@
  • Almost all $\alpha$: A property that holds for all $\alpha$ in a subset of $[0,1]$ with the Lebesgue measure $1$.
  • $\mathds{1}_{E}(*)$: The indicator function of $E$, that is, $\mathds{1}_{E}(x)$ is equal to 1 if $x\in E$ and 0 otherwise. We adopt the convention that $\mathds{1}(x=0)$ denotes $\mathds{1}_{\{0\}}(x)$.
@@ 2.2. General lower and upper bounds
Our first main result provides nearly optimal lower and upper bounds for the minimal gap $\delta_{\min}^{\alpha}(N)$ for sequences with a positive uniform separation. The limits are given in terms of the quantity $D_{N}$, the number of distinct positive differences between the first $N$ terms.
@@
In particular, condition $\inf_{m\neq n}|a_{m}-a_{n}|\geq 1$ is automatically satisfied for any sequence $(a_{n})_{n\geq 1}$ of distinct positive integers. Therefore, we recover the second part of [ABM, Theorem 1].
@@ 2.5. A zero-one law for the minimal gap
Let $\left(a_{n}\right)_{n\geq 1}$ be a sequence of positive real numbers such that $\inf_{1\leq m\neq n\leq N}|a_{m}-a_{n}|\geq 1$. Let $(\eta(N))_{N\geq 1}$ be a sequence of non-negative reals. Let $\mathcal{A}$ denote the set of those $\alpha\in[0,1]$ for which $\lfloor\delta_{\min}^{\alpha}\rfloor(N)\leq\eta(N)$ holds for infinitely many $N$. Then, we have $\lambda(\mathcal{A})=0$ or $\lambda(\mathcal{A})=1$, depending on whether the series
@@
Theorem 13 provides a sharp criterion reducing the problem of determining the almost sure behaviour of the minimal gap of $(\alpha a_{n})_{n\geq 1}$ to the arithmetic structure of $\lfloor(A_{N}-A_{N})^{+}\rfloor$, showing that this floored difference set alone determines the typical asymptotic order, at least in principle.
@@ 2.6. Applications and examples
  • Lacunary sequences: Sequences that satisfy $\frac{a_{n+1}}{a_{n}}\geq D>1$, such as $a_{n}=e^{n}$ or $\pi^{n}$.
@@
Now, taking the minimum over all $m^{\prime}\neq n^{\prime}\leq N$, we obtain $\widetilde{\delta}_{\min}^{\alpha}(N)\leq\widehat{\delta}_{\min}^{\alpha}(N)$.
@@
Now, taking the minimum over all $m^{\prime}\neq n^{\prime}\leq N$, we obtain $\delta_{\min}^{\alpha}(N)\leq\widetilde{\delta}_{\min}^{\alpha}(N)$. Therefore, for all $N\in\mathbb{N}$, we obtain